OpenStudy (anonymous):

Solve the equation for the letter y: -2x + 4y = 8 i know how to do it i just wanna know if this is the right way to do it ???? -2x + 4y = 8 4y=8-2x y=(8-2x)4 or y=(8-2x)/4

OpenStudy (anonymous):

-2x + 4y = 8 4y=8+2x y=(8+2x)/4

OpenStudy (anonymous):

and u are sure about this

OpenStudy (anonymous):

@johana -2x + 4y = 8 4y=8-2x it should be 4y = 8 + 2x not -2x that's the only difference between your answers

OpenStudy (campbell_st):

start by adding 2 x to both sides what do you get...?

OpenStudy (anonymous):

For the equation. -2x + 4y = 8 y = (x + 4)/2